OpenStudy (blueyoshi):

WILL FAN AND MEDAL PLEASE HELP QUICKLY line EC bisects angle BED, m angle AEB = 11x-12 and m angle CED = 4x + 1. Find m angle AEC

OpenStudy (blueyoshi):

@imnotarobot5 @paigephay_31903 @_Malice_ @butlertechrules

OpenStudy (blueyoshi):

Sorry I'm like freaking the heck out, I'm just in a big hurry...

OpenStudy (imnotarobot5):

When you say m angle BED what do you mean?

OpenStudy (blueyoshi):

I mean that's the name of the angle, the problem has an m then an angle symbol and then BED

OpenStudy (imnotarobot5):

Is that the full question? Where is A supposed to be?

OpenStudy (blueyoshi):

That's the picture of the angles

OpenStudy (imnotarobot5):

Alright so line AD= 180 because any line is 180 right?

OpenStudy (blueyoshi):

right

OpenStudy (blueyoshi):

(btw just ignore the "m's" in the problem I typed up

OpenStudy (imnotarobot5):

so from that you can retricethe equations of angles BEC CED and AEB together

OpenStudy (imnotarobot5):

Because it gives you CED as 4x+1 and it tells you that BED is bisected, that means CED is equal to BEC

OpenStudy (imnotarobot5):

so 4x+1 + 4x+1 + 11x - 12

OpenStudy (imnotarobot5):

then set that all equal to 180 because that whole side of the line equals 180

OpenStudy (blueyoshi):

okay

OpenStudy (imnotarobot5):

19x-10 = 180 add 10 to both sides 19x = 190 divide by 19 on both sides x = 10

OpenStudy (imnotarobot5):

now that you know x, you can add the equation of AEB(which is 11x-12) and the equation of BEC(which is 4x+1) together so 11x-12+4x+1 15x +11 then just insert the 10 you found 15(10) + 11 = 150 +11 161

OpenStudy (blueyoshi):

THANK YOU SO MUCH!!

OpenStudy (imnotarobot5):

sorry, the answer is 139

OpenStudy (imnotarobot5):

Go with 139! Just subtract 4(10) + 1 from 180

OpenStudy (imnotarobot5):

= 139

OpenStudy (blueyoshi):

oh okay

OpenStudy (blueyoshi):

Thank you again!!